OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12

OpenStudy (anonymous):

-11x+ 9y = -12 add 11x to both sides +11x +11x __________________ 9y = 11x -12 divide 9 to both sides y = \[\frac{ 11 }{ 9 }x -\frac{ 12 }{ 9 }\] You can simplify \[-\frac{ 12 }{ 9}\]to \[-\frac{ 4 }{ 3}\] so y= \[\frac{ 11 }{ 9 } x - \frac{ 4 }{ 3}\] and the slope in the y=mx+b form is the "m" part , Rise over run, the y -intercept is the "b"

OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4)

OpenStudy (anonymous):

wait but thats not a answer

OpenStudy (anonymous):

nvm!! sorry thankssss

OpenStudy (anonymous):

Really? Hmm.. well the slope would be 11/9x and Y-intercept is -4/3 , or I must be wrong somewhere. and use the slope formula on (6,2) and (7,4) x\[x _{1} = 6 , x _{2} = 2 , y _{1} = 7 , y _{2} =4\] m = \[y _{2} - y _{1} / x _{2} - x _{1}\]

OpenStudy (anonymous):

Just plug it in and you'll find the slope